A decrease in the price of a product from R40,00 to R30,00 causes the quantity demanded to increase from 2 000 to 2 500 units. Using the arc elasticity of demand, the price elasticity of demand is


1
Expert's answer
2021-06-15T11:54:13-0400

Arc Elasticity of Demand="\Delta{Q} \over\Delta{P}" "\times{P1 +P2}\over{Q1+Q2}"

="{2500-2000}\over{30000-40000}" "\times" "{40000+30000}\over2000+2500"


="\ 500\over -10000" "\times" "\ 70000\over 4500"

Price Elasticity Of Demand ="\ 0.78"